The plane making intercept 1 at the z-axis and parallel to the xy- plane intersects the cone x^2+y^2=z^2tan^2 ϑ in a circle.
True or false with full explanation
1
Expert's answer
2020-12-08T06:40:46-0500

The equation of the circle with center "O(a,b)" and radius "R" is "(x-a)^2+(y-b)^2=R^2". if "z=1" then the equation of the intersection of the "xy"- plane and the cone "x^2+y^2=z^2tan^2 \u03d1" is "x^2+y^2=\tan^2 ϑ". So, it is a circle with center "O(0,0,1)" and radius "R=\tan ϑ" containing in the plane "z=1".


Answer: true